OpenStudy (cheesecakekitten):

WILL GIVE MEDAL AND FAN!!!!!!!!!!!!!!!! Ellies bird feeder is shaped like a cone with a height of 25 in. and a radius of 3 in. Packages of bird seed are cylindrical and come in different sizes. What is the height of the package that has a radius of 5 in. and contains exactly enough bird seed to fill the feeder? Use 3.14 to approximate pi.

OpenStudy (cheesecakekitten):

@iGreen

OpenStudy (cheesecakekitten):

Are you people going to help me or stare

OpenStudy (igreen):

Stare

OpenStudy (igreen):

xD

OpenStudy (cheesecakekitten):

.-.

OpenStudy (igreen):

Hold on a bit..

OpenStudy (cheesecakekitten):

kk

OpenStudy (igreen):

Okay, let's find the volume of the cone first.. \(V = \dfrac{1}{3} \pi r^2 h\) Plug in what we know: \(V = \dfrac{1}{3} (3.14) (3)^2 (25)\) Simplify that

OpenStudy (cheesecakekitten):

on it

OpenStudy (cheesecakekitten):

235.5

OpenStudy (igreen):

Yep.

OpenStudy (cheesecakekitten):

so that's the answer?

OpenStudy (igreen):

No..hold on

OpenStudy (cheesecakekitten):

it should be the heigh just saying

OpenStudy (igreen):

Let's see, the volume of a cone is 1/3 the volume of a cylinder.

OpenStudy (cheesecakekitten):

*height

OpenStudy (igreen):

Volume of a cylinder formula: \(V = \pi r^2 h\) Plug in 235.5 for the volume and 5 for the radius, and solve for 'h': \(235.5 = (3.14)(5)^2 h\) Multiply 25 * 3.14

OpenStudy (cheesecakekitten):

78.5

OpenStudy (igreen):

Yes, so we have: \(235.5 = 78.5h\) Divide 78.5 to both siddes

OpenStudy (igreen):

*sides

OpenStudy (cheesecakekitten):

like, 235.5 / 78.5 and 78.5/78.5?

OpenStudy (igreen):

Yep

OpenStudy (cheesecakekitten):

kk

OpenStudy (cheesecakekitten):

235.5 / 78.5 = 3 and 78.5/78.5 = 1

OpenStudy (igreen):

Yes, so we have: \(h = 3\) So the height is 3, that's your answer.

OpenStudy (cheesecakekitten):

^_^ thnx so much

OpenStudy (igreen):

Np

OpenStudy (cheesecakekitten):

i have another